### v1

# Structural Nested Mean Models for Modified Treatment Policies

Abstract
There is a growing literature on estimating effects of treatment strategies based on the natural treatment that would have been received in the absence of intervention, often dubbed ‘modified treatment policies’ (MTPs). MTPs are sometimes of interest because they are more realistic than interventions setting exposure to an ideal level for all members of a population. In the general time-varying setting, Richardson and Robins [2013] provided exchangeability conditions for nonparametric identification of MTP effects that could be deduced from Single World Intervention Graphs (SWIGs). Díaz et al. [2023] provided multiply robust estimators under these identification assumptions that allow for machine learning nuisance regressions. In this paper, we fill a remaining gap by extending Structural Nested Mean Models (SNMMs) [Robins, 1994a, 2004, Vansteelandt and Joffe, 2014] to MTP settings, which enables characterization of (time-varying) heterogeneity of MTP effects. We do this both under the exchangeability assumptions of Richardson and Robins [2013] and under parallel trends assumptions, which enables investigation of (time-varying heterogeneous) MTP effects in the presence of some unobserved confounding.
1 Introduction
There is a growing literature on estimating effects of treatment strategies based on the natural treatment that would have been received in the absence of intervention [Robins et al., 2004, Richardson and Robins, 2013, Young et al., 2014, Haneuse and Rotnitzky, 2013, Muñoz and Van Der Laan, 2012, Díaz et al., 2023, Sani et al., 2020]. Examples of such strategies include ‘exercise 20 minutes longer than you normally would’ or ‘discharge patients from the intensive care unit one day later than under usual care’. Strategies depending on the natural value of treatment have been dubbed ‘modified treatment policies’ (MTPs), a term which has gained some traction and we hence adopt here. MTPs are sometimes of interest because they are more realistic than interventions setting exposure to an ideal level for all members of a population.
In the general time-varying setting, Richardson and Robins [2013] provided exchangeability conditions for nonparametric identification of MTP effects that could be deduced from Single World Intervention Graphs (SWIGs). Díaz et al. [2023] provided multiply robust estimators under these identification assumptions that allow for machine learning nuisance regressions. In this paper, we fill a remaining gap by extending Structural Nested Mean Models (SNMMs) [Robins, 1994a, 2004, Vansteelandt and Joffe, 2014] to MTP settings, which enables characterization of (time-varying) heterogeneity of MTP effects. We do this both under the exchangeability assumptions of Richardson and Robins [2013] and under parallel trends assumptions [Shahn et al., 2022], which enables investigation of (time-varying heterogeneous) MTP effects in the presence of some unobserved confounding.
The structure of the paper is as follows. In Section 2, we summarize notation and review definitions of MTPs and SNMMs. Once defined, we also provide some hypothetical examples of when SNMMs for MTP effects might be of interest. In Section 3, we provide identification results and Neyman orthogonal estimators under exchangeability assumptions. In Section 4, we provide identification results under parallel trends assumptions. In Section 5, we provide simulations showing that our estimators are unbiased and that the sandwich variance estimates provide nominal coverage. In Section 6, we present a real data analysis estimating the effects of shifting mobility from its natural level on subsequent county level Covid-19 incidence. In Section 7, we conclude and discuss some directions for future work.
2 Notation, MTPs, and SNMMs
Suppose we observe a cohort of $N$ subjects indexed by $i\in\{1,\ldots,N\}$. Assume that each subject is observed at regular intervals from baseline time 0 through end of follow-up time $K$, and there is no loss to follow-up. At each time point $t$, the data are collected on $O_{t}=(Z_{t},Y_{t},A_{t})$ in that temporal order. $A_{t}$ denotes the (possibly multidimensional with discrete and/or continuous components) treatment received at time $t$, $Y_{t}$ denotes the outcome of interest at time $t$, and $Z_{t}$ denotes a vector of covariates at time $t$ excluding $Y_{t}$. Hence, $Z_{0}$ constitutes the vector of baseline covariates other than $Y_{0}$. For arbitrary time varying variable $X$: we denote by $\bar{X}_{t}=(X_{0},\ldots,X_{t})$ the history of $X$ through time $t$; we denote by $\underline{X}_{t}=(X_{t},\ldots,X_{K})$ the future of $X$ from time $t$ through time $K$; and whenever the negative index $X_{-1}$ appears it denotes the null value with probability 1. In Section 3, when we work under exchangeability assumptions, we define $\bar{L}_{t}$ to be $(\bar{Z}_{t},\bar{Y}_{t})$, i.e. the joint covariate and outcome history through time $t$, but we do not require that the outcome is necessarily measured at all time points. In Section 4, when we work under parallel trends assumptions, we define $\bar{L}_{t}$ to be $(\bar{Z}_{t},\bar{Y}_{t-1})$, i.e. the joint covariate and outcome history through time $t$ excluding the most recent outcome $Y_{t}$, and we assume that the outcome is measured at each time point. Hence, in the parallel trends section, $L_{0}$ is $Z_{0}$. In all sections, we let $H_{t}=(\bar{L}_{t},\bar{A}_{t-1})$ denote the relevant pre-treatment history. We denote random variables by capital letters and their realizations using lower case letters. We adopt the counterfactual framework for time-varying treatments [Robins, 1986] which posits that corresponding to each time-varying treatment regime $\bar{a}_{t}$, each subject has a counterfactual or potential outcome $Y_{t+1}(\bar{a}_{t})$ that would have been observed had that subject received treatment regime $\bar{a}_{t}$.
Let $g=(g_{1},\ldots,g_{K})$ denote an arbitrary MTP with each $g_{t}:(h_{t},a_{t})\rightarrow a_{t}^{+}$ a treatment rule setting the modified treatment ($a_{t}^{+}$) as a function of observed history through $t$. $A_{t}(g)$ is the treatment that would naturally occur at time $t$ had strategy $g$ been imposed through t1t-1. $A_{t}^{+}(g)$ is the treatment received at time $t$ under $g$, possibly a function of $A_{t}(g)$. $Y_{t}(g)\equiv Y_{t}(A_{t-1}^{+}(g))$ is the counterfactual outcome had treatment been assigned according to $g$. We will use the shorthand $Y_{t}(\bar{A}_{m},\underline{g})$ to denote the counterfactual outcome under the observed regime through time $m$ followed by regime $g$ thereafter.
A general regime SNMM for MTP $g$ models the contrasts
$\gamma_{tk}^{g}(h_{t},a_{t})=E[Y_{k}(\bar{A}_{t},\underline{g})-Y_{k}(\bar{A}_{t-1},\underline{g})|H_{t}=h_{t},A_{t}=a_{t}].$ (1)
(1) represents the conditional lasting effects among subjects with observed history $(h_{t},a_{t})$ of receiving the observed treatment at time $t$ then switching to MTP $g$ thereafter, compared to switching to $g$ at time $t$ and continuing to follow it thereafter. Thus, (1) might be interpreted as the conditional effect of one final blip of the observational regime before switching to an MTP.
With knowledge of $\gamma_{tk}^{g}$, we can ‘blip down’ or strip away these effects from observed outcomes to obtain consistent estimates of conditional counterfactual outcomes under the MTP $g$.
Lemma 1.
$E[Y_{k}-\sum_{j=t}^{k-1}\gamma_{j}(H_{j},A_{j})|H_{t},A_{t}]=E[Y_{k}(\bar{A}_{t-1},\underline{g})|H_{t},A_{t}]$
The proof of Lemma 1 is in the Appendix. It follows as a special case of Lemma 1 that given $\gamma_{tk}^{g}$, $E[Y_{k}(g)]$ is identified as $E[Y_{k}-\sum_{j=0}^{k-1}\gamma_{jk}(A_{j},H_{j})]$.
A parametric SNMM specifies a functional form
$\gamma_{tk}^{g}(h_{t},a_{t})=\gamma_{tk}^{g}(h_{t},a_{t};\psi^{*})$ (2)
with $\gamma_{tk}^{g}(h_{t},a_{t};\psi)$ a known function of finite dimensional parameter $\psi$ taking unknown true value $\psi^{*}$ such that $\gamma_{tk}^{g}(h_{t},a_{t};\psi)=0$ whenever $g(h_{t},a_{t})=a_{t}$ or $\psi=0$.
As one substantive example, suppose $A_{t}$ is a continuous measure of medication adherence (e.g. proportion of prescribed doses taken), and it is thought that efficacy starts significantly declining below some threshold $\delta$. One might be interested in the effect of a partial adherence enforcing intervention $g(h_{t},a_{t})=\mathbf{1}\{a_{t}<\delta\}\delta+\mathbf{1}\{a_{t}\geq\delta\}a_{t}$ that sets treatment to $\delta$ if its natural value is below that threshold and leaves the natural value unchanged otherwise. In this setting, $\gamma_{tk}(h_{t},a_{t})$ is the lasting effect of adherence level $a_{t}$ versus $g(h_{t},a_{t})$ in month $t$ followed by regime $g$ thereafter. If $a_{t}\geq\delta$, then the effect would be 0. For $a_{t}<\delta$, parametric model (2) could characterize how the lasting effects of a blip of non-adherence relative to regime $g$ depend on the magnitude of nonadherence ($\delta-a_{t}$) and health history.
As another example, this time concerning an exposure more often studied under parallel trends assumptions, suppose $A_{t}$ denotes state minimum wage in year $t$. Perhaps interest centers on the effect on poverty of increasing minimum wage by $\mathdollar 2$ above its natural value if its natural value is less than $\mathdollar 10$. (1) then defines the conditional effects on poverty of natural wages less than $\mathdollar 10$ relative to their value had they been increased by $\mathdollar 2$. (1) might, for example, reveal that $\mathdollar 2$ wage hikes relative to observed wages would have been more impactful in states that actually had lower wages. This is a different estimand than the effect of treatment on the treated (i.e. the effect of wage hikes from the previous year in states that implemented them) that economists typically study with Difference in Differences.
3 Identification and Estimation Under Exchangeability Assumptions
3.1 Point Exposure Setting
In the point exposure setting where $K=1$, we will drop time subscripts for simplicity. We will make the standard consistency assumption that
$\textbf{Consistency: }Y(a)=Y\text{ whenever }A=a.$ (3)
We also make a positivity assumption with respect to regime $g$
$f_{A,L}(a,l)>0\text{ implies }f_{A,L}(g(a,l),l)>0,$ (4)
which states that for any treatment and covariate values that might be observed in the data, the corresponding $g$-modified treatment value with the same covariates might also be observed. Finally, Haneuse and Rotnitzky [2013] introduced the MTP exchangeability assumption
$E[Y(a^{\prime})|A=a,L=l]=E[Y(a^{\prime})|A=a^{\prime},L=l]\text{ for all }a^{\prime}=g(a,l)$ (5)
and showed that under assumptions (3), (4), and (5), conditional MTP effects are identified by
$E[Y(g)|A=a,L=l]=E[Y|A=g(a,l),L=l].$ (6)
Identification of a SMM follows immediately:
$\gamma^{g}(l,a)\equiv E[Y-Y(g(l,a))|A=a,L=l]=E[Y|A=a,L=l]-E[Y|A=g(a,l),L=l].$ (7)
Consider a parametric SMM $\gamma^{g}(l,a;\psi^{*})$ as defined in (2). Then (7) implies that $\psi^{*}$ satisfies
$E[q(L,A)(Y-\gamma^{g}(L,A;\psi^{*})-\mu(g(A,L),L))]=0$ (8)
where $\mu(a,l)=E[Y|A=a,L=l]$. Thus, a consistent and asymptotically normal estimator of $\psi^{*}$ can be obtained by solving the corresponding estimating equations
$\mathbb{P}_{N}\begin{bmatrix}b(A,L)\{Y-\mu(A,L;\beta)\}\\ q(A,L)\{Y-\gamma^{g}(A,L;\psi)-\mu(g(A,L);\beta)\}\end{bmatrix}=0$ (9)
where $\mu(A,L;\beta)$ is a model for $\mu(A,L)$ smooth in an $r$-dimensional parameter $\beta$ (equal to $\beta^{*}$ under the true law) and $q(A,L)$ and $b(A,L)$ are conformable analyst selected index functions.
The estimator solving (9) is very sensitive to misspecification of $\mu$, and asymptotic normality may not hold if, to avoid misspecification, $\mu$ is estimated by machine learning. Therefore, we introduce an estimator that has the Neyman-orthogonality property [Chernozhukov et al., 2018] crucial for enabling machine learning estimation of nuisance functions.
Let $T_{g}$ act on functions $h$ by $(T_{g}h)(A,L)=h(g(A,L),L)$, and define the $L$-conditional inner product weighted by the observed treatment density:
$\langle u,v\rangle\;=\;E\!\left[\int u(L,a)\,v(L,a)\,f_{A\mid L}(a\mid L)\,da\right].$
Let $T_{g}^{\dagger}$ be the adjoint to $T_{g}$ such that
$\langle q,T_{g}h\rangle_{w}=\langle T_{g}^{\dagger}q,h\rangle_{w}.$
For a shift $g(a)=a+\delta$,
$(T_{g})^{\dagger}q(L,A)\;=\;q\!\big(L,g(A)\big)\,\frac{f_{A\mid L}(A-\delta\mid L)}{f_{A\mid L}(A\mid L)}.$
In particular, for $\tilde{q}:=T_{g}^{\dagger}q$ we have the identity
$E[q(A,L)\,h\!\big(g(A,L),L\big)\,f_{A|L}(A|L)]=E\!\left[\tilde{q}(A,L)\,h(A,L)\,f_{A|L}(A|L)\right]\quad\forall h.$ (10)
Thus, $T_{g}^{\dagger}$ is essentially a change of variables operator for $g$. If $g(\cdot,l)$ is bijective and differentiable, then
$\tilde{q}(l,a)=q\!\big(l,g^{-1}(a,l)\big)\,\frac{\pi\!\big(g^{-1}(a,l)\mid l\big)}{\pi(a\mid l)}\,\big|\det J_{a}g^{-1}(a,l)\big|.$
If $g(\cdot,l)$ is possibly many-to-one and $A$ is discrete,
$\tilde{q}(l,a)=\sum_{a^{\prime}:\,g(a^{\prime},l)=a}q(l,a^{\prime})\,\frac{\pi(a^{\prime}\mid l)}{\pi(a\mid l)}.$
Theorem 1.
(Neyman orthogonal estimator, point exposure) Assume Consistency, Positivity, MTP Exchangeability, and regularity conditions. Consider the score
$\phi(O;\psi,\eta)=\underbrace{(q-\tilde{q})\,(Y-\mu(A,L))}_{\text{augmentation}}+\underbrace{q\{\mu(A,L)-\mu(g(A,L),L)-\gamma_{g}(L,A;\psi)\}}_{\text{identifying}},\quad\eta=(\mu,\pi).$ (11)
Then:
  1. 1.
    Identification. At $(\psi^{\ast},\eta^{\ast})$, $E[\phi(O;\psi^{\ast},\eta^{\ast})]=0.$
  2. 2.
    Neyman orthogonality. For any regular parametric submodel $t\mapsto\eta_{t}=(\mu_{t},\pi_{t})$ with $\eta_{0}=\eta^{\ast}$,
    $\left.\frac{d}{dt}\right|_{t=0}E\!\left[\phi(O;\psi^{\ast},\eta_{t})\right]=0.$
  3. 3.
    Asymptotic variance Assume that $\hat{\eta}$ is learned on held-out folds and $\hat{\eta}-\eta^{*}||_{L_{2}}=o_{p}(1)$. If $\dim(q)=\dim(\psi)=d$ and $\hat{\psi}$ solves $\frac{1}{n}\sum_{i=1}^{n}\hat{\phi}_{i}(\psi)=0$, then for
    $G:=E[\frac{\partial}{\partial\psi^{\top}}\phi(O;\psi^{*},\eta^{*})]\in R^{d\times d}$
    and
    $\Sigma:=Var(\phi(O;\psi^{*},\eta^{*})\big)\in R^{d\times d},$
    $\sqrt{n}(\hat{\psi}-\psi^{*})\;\rightsquigarrow\;\mathcal{N}\!\big(0,\;V\big),\qquad IF(O)=-G^{-1}\phi(O;\psi^{*},\eta^{*}),\quad V=G^{-1}\Sigma(G^{-1})^{\top}.$
Proof.
See Appendix 8.1. ∎
Algorithm 1 computes the estimator from Theorem 1.
Algorithm 1 Cross-fitted Neyman–orthogonal estimator for a point-exposure MTP
1:Data $\{(Y_{i},A_{i},H_{i})\}_{i=1}^{n}$; basis $s(H)$; shift $\delta$; folds $K$
[…]
3:Partition indices into $K$ folds $(\mathcal{T}^{(k)},\mathcal{I}^{(k)})$
4:Initialize $M\leftarrow 0$, $b\leftarrow 0$
5:for $k=1,\dots,K$ do $\triangleright$ Nuisance fits on train fold
6:  Fit $\hat{\mu}$ on $\{(Y_{j},A_{j},H_{j})\}_{j\in\mathcal{T}^{(k)}}$ for $y\approx E[Y\mid A,H]$
7:  Fit $\hat{m}$ on $\{(A_{j},H_{j})\}_{j\in\mathcal{T}^{(k)}}$ for $a\approx E[A\mid H]$
8:  Estimate $\hat{\sigma}$ as SD of residuals $A-\hat{m}(H)$ on $\mathcal{T}^{(k)}$ (if using Normal ratio)
9:  for all $i\in\mathcal{I}^{(k)}$ do
10:   $e_{i}\leftarrow Y_{i}-\hat{\mu}(A_{i},H_{i})$; $\mu^{g}_{i}\leftarrow\hat{\mu}(A_{i}+\delta,H_{i})$
11:   $s_{i}\leftarrow s(H_{i})$; $q_{i}\leftarrow s_{i}$
12:   if Normal working residual for $A\!\mid\!H$ then
13:     $\hat{r}_{i}\leftarrow\dfrac{\phi(A_{i}-\delta;\ \hat{m}(H_{i}),\hat{\sigma})}{\phi(A_{i};\ \hat{m}(H_{i}),\hat{\sigma})}$
14:   else
15:     Estimate $\hat{r}_{i}\approx\dfrac{\hat{f}_{A\mid H}(A_{i}-\delta\mid H_{i})}{\hat{f}_{A\mid H}(A_{i}\mid H_{i})}$
16:   $\tilde{q}_{i}\leftarrow q_{i}\cdot\hat{r}_{i}$ $\triangleright$ Accumulate normal equations $M\hat{\psi}=b$
17:   $M\leftarrow M+\delta\,s_{i}s_{i}^{\top}$
18:   $b\leftarrow b-\big[s_{i}\{\hat{\mu}(A_{i},H_{i})-\mu^{g}_{i}\}+(q_{i}-\tilde{q}_{i})\,e_{i}\big]$
19:Solve and variance:
[…]
3.2 Time varying treatments
For didactic purposes, consider identification in the two time point case. We extend the results by induction to the general time varying case in the Appendix. We know that the one time step ahead effects $\gamma_{12}^{g}(a_{1},h_{1})$ and $\gamma_{01}^{g}(a_{0},l_{0})$ are identified from the point exposure case. So we will focus on identification of $\gamma_{02}^{g}(a_{0},l_{0})$.
$\gamma_{02}^{g}(a_{0},l_{0})=E[Y_{2}(a_{0},g_{1})-Y_{2}(g_{0},g_{1})|L_{0}=l_{0},A_{0}=a_{0}]$
By iterated expectations and Lemma 1,
$E[Y_{2}(a_{0},g_{1})|A_{0}=a_{0},L_{0}=l_{0}]$
$=E[E[Y_{2}(a_{0},g_{1})|A_{0}=a_{0},L_{0}=l_{0},L_{1},A_{1}]|A_{0}=a_{0},L_{0}=l_{0}]$
$=E[E[Y_{2}-\gamma_{12}^{g}(A_{1},H_{1})|A_{0}=a_{0},L_{0}=l_{0},L_{1},A_{1}]|A_{0}=a_{0},L_{0}=l_{0}]$
$=E[Y_{2}-\gamma_{12}^{g}(A_{1},H_{1})|A_{0}=a_{0},L_{0}=l_{0}]$
Thus, the first term of $\gamma_{02}^{g}(a_{0},l_{0})$ is identified. The second term is identified as
$E[Y_{2}(g_{0},g_{1})|L_{0}=l_{0},A_{0}=a_{0}]=$
$\sum_{l_{1},a_{1}}E[Y_{2}|L_{0}=l_{0},A_{0}=g(l_{0},a_{0}),L_{1}=l_{1},A_{1}=g(l_{0},g(l_{0},a_{0}),a_{1})]$
$p(l_{1}|L_{0}=l_{0},A_{0}=g(l_{0},a_{0}))p(a_{1}|L_{0}=l_{0},L_{1}=l_{1},A_{0}=g(l_{0},a_{0}))$
$\equiv\mu_{02}(l_{0},g(l_{0},a_{0}))$
by the extended g-formula under the assumptions of Section 5 of [Richardson and Robins, 2013]. Following Díaz et al. [2023], $\mu_{02}(l_{0},a_{0})$ can be represented in terms of iterative regressions by
$\mu_{02}(l_{0},a_{0})=E[\mu_{12}(l_{0},L_{1},g(l_{0},a_{0}),g(l_{0},L_{1},g(l_{0},a_{0}),A_{1}))|L_{0}=l_{0},A_{0}=a_{0}]$ (12)
with $\mu_{12}(l_{0},l_{1},a_{0},a_{1})=E[Y_{2}|L_{0}=l_{0},L_{1}=l_{1},A_{0}=a_{0},A_{1}=a_{1}]$. Then for parametric model $(\gamma_{01}^{g}(l_{0},a_{0}),\gamma_{12}^{g}(h_{1},a_{1}),\gamma_{02}^{g}(l_{0},a_{0}))=(\gamma_{01}^{g}(l_{0},a_{0};\psi^{*}),\gamma_{12}^{g}(h_{1},a_{1};\psi^{*}),\gamma_{02}^{g}(l_{0},a_{0};\psi^{*}))$, parameter $\psi^{*}$ satisfies
$E\begin{bmatrix}q_{01}(L_{0},A_{0})\{Y_{1}-\gamma_{01}^{g}(L_{0},A_{0};\psi^{*})-\mu_{01}(L_{0},g(L_{0},A_{0}))\}\\ q_{12}(H_{1},A_{1})\{Y_{1}-\gamma_{12}^{g}(H_{1},A_{1};\psi^{*})-\mu_{12}(H_{1},g(H_{1},A_{1}))\}\\ q_{02}(L_{0},A_{0})\{Y_{1}-\gamma_{12}^{g}(H_{1},A_{1};\psi^{*})-\gamma_{02}^{g}(L_{0},A_{0};\psi^{*})-\mu_{02}(L_{0},g(L_{0},A_{0}))\}\end{bmatrix}=0,$ (13)
which again leads to natural estimating equations. In particular, a consistent and asymptotically normal estimator of $\psi^{*}$ can be obtained by solving
$\mathbb{P}_{N}\begin{bmatrix}b_{01}(L_{0},A_{0})\{Y_{1}-\mu_{01}(L_{0},A_{0};\beta)\\ b_{12}(\bar{L}_{1},\bar{A}_{1})\{Y_{2}-\mu_{12}(\bar{L}_{1},\bar{A}_{1};\beta)\\ b_{02}(L_{0},A_{0})\{\mu_{12}(L_{0},L_{1},g(L_{0},A_{0}),g(L_{0},L_{1},g(L_{0},A_{0}),A_{1}))-\mu_{02}(L_{0},A_{0};\beta)\}\\ q_{01}(L_{0},A_{0})\{Y_{1}-\gamma_{01}^{g}(L_{0},A_{0};\psi)-\mu_{01}(L_{0},g(L_{0},A_{0}))\}\\ q_{12}(H_{1},A_{1})\{Y_{1}-\gamma_{12}^{g}(H_{1},A_{1};\psi)-\mu_{12}(H_{1},g(H_{1},A_{1}))\}\\ q_{02}(L_{0},A_{0})\{Y_{1}-\gamma_{12}^{g}(H_{1},A_{1};\psi)-\gamma_{02}^{g}(L_{0},A_{0};\psi)-\mu_{02}(L_{0},g(L_{0},A_{0}))\}\end{bmatrix}=0$ (14)
where $\mu_{mk}(\bar{A}_{m},\bar{L}_{m};\beta)$ are models for $\mu_{mk}(\bar{A}_{m},\bar{L}_{m})$ smooth in an $r$-dimensional parameter $\beta$ (equal to $\beta^{*}$ under the true law) and $q_{mk}(\bar{A}_{m},\bar{L}_{m})$ and $b_{mk}(\bar{A}_{m},\bar{L}_{m})$ are conformable analyst selected index functions.
As in the point exposure setting, a Neyman orthogonal estimator would be desirable to enable machine learning estimation of nuisance functions. We now develop this estimator for the general time-varying setting. Define the one step ahead conditional mean
$\mu_{t}(a,h)\equiv E[V_{t+1}|A_{t}=a,H_{t}=h]$
where we set $V_{T}=Y$ and recursively set for $t=T-1,\ldots,0$
$V_{t}=\mu_{t}(g(H_{t},A_{t}),H_{t}).$
Now, for each time $t$, define the operator $T_{gt}h\equiv h(H_{t},g(H_{t},A_{t}))$. And let $T_{gt}^{\dagger}$ be its adjoint with respect to
$\langle u,v\rangle_{w_{t}}\equiv E[u(H_{t},A_{t})v(H_{t},A_{t})w_{t}]$
for $w_{t}=\pi(A_{t}|H_{t})$. As in the point exposure setting, let $\tilde{q}_{t}\equiv T_{gt}^{\dagger}q_{t}$.
Theorem 2 (Neyman orthogonal estimator, time-varying).
Assume Sequential consistency, positivity, and MTP exchangeability hold (i.e. the assumptions of Section 5 of Richardson and Robins [2013]). Consider the score
$\Phi(O;\psi,\eta)=\sum_{t=0}^{T-1}\phi_{t}(O;\psi,\eta),$
with
$\phi_{t}=\big(q_{t}-\tilde{q}_{t}\big)\,\{V_{t+1}-\mu_{t}(H_{t},A_{t})\}+q_{t}\Big\{\mu_{t}(H_{t},A_{t})-\mu_{t}(H_{t},g_{t}(A_{t},H_{t}))-\gamma_{t}(H_{t},A_{t};\psi)\Big\},$
where $\eta=\{\mu_{t},\pi_{t},\tilde{q}_{t}:t=0,\dots,T-1\}$.
  1. 1.
    Identification At the truth $(\psi^{\ast},\eta^{\ast})$, $\mathbb{E}[\Phi(O;\psi^{\ast},\eta^{\ast})]=0$
  2. 2.
    Neyman orthogonality The score is Neyman-orthogonal: the pathwise derivative of $\mathbb{E}[\Phi(O;\psi^{\ast},\eta)]$ with respect to each nuisance $\mu_{t}$ and $\pi_{t}$ vanishes at $\eta^{\ast}$.
  3. 3.
    Asymptotic variance Suppose $m=d$ and $\psi$ is estimated by the root $\hat{\psi}$ of the cross-fitted sample moment $n^{-1}\sum_{i=1}^{n}\hat{\phi}_{i}(\psi)=0$, with nuisances learned on held-out folds and consistent in $L_{2}$. Let
    $G:=E\!\left[\frac{\partial}{\partial\psi^{\top}}\phi(O;\psi^{\ast},\eta^{\ast})\right]\in R^{d\times d},\qquad\Sigma:=Var\!\big(\phi(O;\psi^{\ast},\eta^{\ast})\big)\in R^{d\times d}.$
    Then under standard regularity conditions,
    $\sqrt{n}(\hat{\psi}-\psi^{\ast})\;\overset{d}{\to}\;\mathcal{N}\!\big(0,\;V\big),\qquad\text{with IF }\ IF(O)=-\,G^{-1}\phi(O;\psi^{\ast},\eta^{\ast}),$
    and
    $V\;=\;G^{-1}\,\Sigma\,(G^{-1})^{\top}.$
Proof.
See Appendix 8.2
Algorithm 2 implements the estimator from Theorem 2.
Algorithm 2 Cross-fitted Neyman–orthogonal estimator for longitudinal MTPs
1:Data $\{(Y_{i},\{A_{t,i},H_{t,i}\}_{t=0}^{T-1})\}_{i=1}^{n}$; bases $s_{t}(H_{t})$; shifts $\{\delta_{t}\}$; folds $K$
[…]
3:Initialize $V_{T,i}\leftarrow Y_{i}$ for all $i$; $M_{t}\leftarrow 0$, $b_{t}\leftarrow 0$ for $t=0,\dots,T\!-\!1$
4:Partition indices into $K$ folds $(\mathcal{T}^{(k)},\mathcal{I}^{(k)})$
5:for $t=T-1,\dots,0$ do
6:  for $k=1,\dots,K$ do $\triangleright$ Nuisance fits on train fold at time $t$
7:   Fit $\hat{\mu}_{t}$ on $\{(V_{t+1,j},A_{t,j},H_{t,j})\}_{j\in\mathcal{T}^{(k)}}$ for $v\approx E[V_{t+1}\mid A_{t},H_{t}]$
8:   Fit $\hat{m}_{t}$ on $\{(A_{t,j},H_{t,j})\}_{j\in\mathcal{T}^{(k)}}$ for $a\approx E[A_{t}\mid H_{t}]$
9:   Estimate $\hat{\sigma}_{t}$ as SD of $A_{t}-\hat{m}_{t}(H_{t})$ on $\mathcal{T}^{(k)}$ (if using Normal ratio)
10:   for all $i\in\mathcal{I}^{(k)}$ do
11:     $e_{t,i}\leftarrow V_{t+1,i}-\hat{\mu}_{t}(A_{t,i},H_{t,i})$; $\mu^{g}_{t,i}\leftarrow\hat{\mu}_{t}(A_{t,i}+\delta_{t},H_{t,i})$
12:     $s_{t,i}\leftarrow s_{t}(H_{t,i})$; $q_{t,i}\leftarrow s_{t,i}$
13:     if Normal working residual then
14:      $\hat{r}_{t,i}\leftarrow\dfrac{\phi(A_{t,i}-\delta_{t};\ \hat{m}_{t}(H_{t,i}),\hat{\sigma}_{t})}{\phi(A_{t,i};\ \hat{m}_{t}(H_{t,i}),\hat{\sigma}_{t})}$
15:     else
16:      Estimate $\hat{r}_{t,i}\approx\dfrac{\hat{f}_{A_{t}\mid H_{t}}(A_{t,i}-\delta_{t}\mid H_{t,i})}{\hat{f}_{A_{t}\mid H_{t}}(A_{t,i}\mid H_{t,i})}$
17:     $\tilde{q}_{t,i}\leftarrow q_{t,i}\cdot\hat{r}_{t,i}$ $\triangleright$ Accumulate time-$t$ normal equations
18:     $M_{t}\leftarrow M_{t}+\delta_{t}\,s_{t,i}s_{t,i}^{\top}$
19:     $b_{t}\leftarrow b_{t}-\big[s_{t,i}\{\hat{\mu}_{t}(A_{t,i},H_{t,i})-\mu^{g}_{t,i}\}+(q_{t,i}-\tilde{q}_{t,i})\,e_{t,i}\big]$
20:     $V_{t,i}\leftarrow\mu^{g}_{t,i}$ $\triangleright$ backward recursion target
21:Solve and variance:
22:for $t=0,\dots,T-1$ do
23:  $\hat{\psi}_{t}\leftarrow(M_{t}+\lambda_{t}I)^{-1}b_{t}$
[…]
25:For each $i$: $\phi_{t,i}\leftarrow(q_{t,i}-\tilde{q}_{t,i})e_{t,i}+s_{t,i}\{\hat{\mu}_{t}(A_{t,i},H_{t,i})-\hat{\mu}_{t}(A_{t,i}+\delta_{t},H_{t,i})+\delta_{t}\,s_{t,i}^{\top}\hat{\psi}_{t}\}$
26:Stack $\phi_{i}\leftarrow(\phi_{0,i}^{\top},\dots,\phi_{T-1,i}^{\top})^{\top}$; $\mathrm{IF}_{i}\leftarrow G^{-1}\phi_{i}$; $\widehat{V}\leftarrow\mathrm{Var}_{n}(\mathrm{IF}_{i})/n$
4 Identification and Parametric Estimation Under Parallel Trends Assumptions
Recall that in this section $\bar{L}_{t}$ excludes the most recent outcome $Y_{t}$. In the parallel trends setting, we present identification results and the natural accompanying parametric outcome regression based estimators. We leave derivation of Neyman orthogonal estimators under parallel trends for future work.
4.1 Point Exposure Setting
Again, we suppress time subscripts for treatment and the baseline covariate to reduce notational clutter in the point exposure setting. We propose the MTP parallel trends assumption
$E[Y_{1}(g(a,l))-Y_{0}|A=a,L=l]=E[Y_{1}(g(a,l))-Y_{0}|A=g(a,l),L=l]\text{ for all }(a,l)\text{ with }p(a,l)>0.$ (15)
In words, the conditional expected counterfactual trend in the outcome under the MTP does not depend on whether treatment takes its natural or modified value. Under (15), $\gamma^{g}(l,a)$ is identified as follows:
$\gamma^{g}(l,a)=E[Y_{1}-Y_{1}(g(a,l))|A=a,L=l]$
$=E[Y_{1}-Y_{0}|A=a,L=l]-E[Y_{1}(g(a,l))-Y_{0}|A=a,L=l]$
$=E[Y_{1}-Y_{0}|A=a,L=l]-E[Y_{1}(g(a,l))-Y_{0}|A=g(a,l),L=l]$
$=E[Y_{1}-Y_{0}|A=a,L=l]-E[Y_{1}-Y_{0}|A=g(a,l),L=l]$
Thus, MTPs can be studied under parallel trends assumptions. Furthermore, $\psi^{*}$ from (2) satisfies
$E[q(L,A)(Y_{1}-Y_{0}-\gamma^{g}(L,A;\psi^{*})-\mu^{d}(L,g(A,L)))]=0,$ (16)
with $\mu^{d}(l,a)=E[Y_{1}-Y_{0}|L=l,A=a]$, which again suggests a natural estimation procedure. A consistent and asymptotically normal estimator of $\psi^{*}$ can be obtained by solving the corresponding estimating equations
$\mathbb{P}_{N}\begin{bmatrix}b(A,L)\{Y_{1}-Y_{0}-\mu^{d}(L,A;\beta)\}\\ q(A,L)\{Y_{1}-Y_{0}-\gamma^{g}(A,L;\psi)-\mu^{d}(L,A;\beta)\}\end{bmatrix}=0$ (17)
where $\mu^{d}(L,A;\beta)$ is a model for $\mu^{d}(L,A)$ smooth in an $r$-dimensional parameter $\beta$ (equal to $\beta^{*}$ under the true law) and $q(A,L)$ and $b(A,L)$ are conformable analyst selected index functions. Derivation of Neyman orthogonal estimators is left for future work.
4.2 Time-Varying Treatments
In the time-varying setting, we make the parallel trends assumptions
$E[Y_{k}(g)-Y_{k-1}(g)|A_{m}=a_{m},H_{m}=h_{m}]=E[Y_{k}(g)-Y_{k-1}(g)|A_{m}=g(h_{m},a_{m}),H_{m}=h_{m}]\text{ for all }h_{m},a_{m},k>m.$ (18)
In words, conditional expected future counterfactual trends under $g$ from time $m$ onwards do not depend on whether treatment at time $m$ took its natural or modified value. For the two time point setting, these assumptions can be enumerated
$\begin{split}E[Y_{2}(g)-Y_{1}(g(a_{0},l_{0}))|A_{0}=a_{0},L_{0}=l_{0}]=E[Y_{2}(g)-Y_{1}(g(l_{0},a_{0}))|A_{0}=g(l_{0},a_{0}),L_{0}=l_{0}]\\ E[Y_{1}(g(l_{0},a_{0}))-Y_{0}|A_{0}=a_{0},L_{0}=l_{0}]=E[Y_{1}(g(l_{0},a_{0}))-Y_{0}|A_{0}=g(l_{0},a_{0}),L_{0}=l_{0}]\\ E[Y_{2}(g)-Y_{1}|A_{1}=a_{1},H_{1}=h_{1}]=E[Y_{2}(g)-Y_{1}|A_{1}=g(h_{1},a_{1}),H_{1}=h_{1}].\end{split}$ (19)
We know that the one time step ahead effects $\gamma_{12}^{g}(a_{1},h_{1})$ and $\gamma_{01}^{g}(a_{0},l_{0})$ are identified from the point exposure case. Identification of $\gamma_{02}^{g}(a_{0},l_{0})$ can be demonstrated as follows.
$\gamma_{02}(a_{0},l_{0})=E[Y_{2}(a_{0},g)-Y_{2}(g)|A_{0}=a_{0},L_{0}=l_{0}]$
$=E[(Y_{2}(a_{0},g)-Y_{1}(g))-(Y_{2}(g)-Y_{1}(g))|A_{0}=a_{0},L_{0}=l_{0}]$
The first term of the conditional expectation on the right hand side is identified as
$E[Y_{2}(a_{0},g)-Y_{1}(g)|A_{0}=a_{0},L_{0}=l_{0}]=E[Y_{2}-\gamma_{12}(l_{0},L_{1},a_{0},A_{1})-(Y_{1}-\gamma_{01}(l_{0},a_{0}))|A_{0}=a_{0},L_{0}=l_{0}].$
The second term is identified as
$E[Y_{2}(g)-Y_{1}(g)|A_{0}=a_{0},L_{0}=l_{0}]$
$=E[Y_{2}(g)-Y_{1}(g)|A_{0}=g(a_{0},l_{0}),L_{0}=l_{0}]$
$=\int_{l_{1},a_{1}}E[Y_{2}(g)-Y_{1}|A_{0}=g(a_{0},l_{0}),A_{1}=a_{1},L_{0}=l_{0},L_{1}=l_{1}]p(l_{1},a_{1}|A_{0}=g(a_{0},l_{0}),L_{0}=l_{0})da_{1}dl_{1}$
$=\int_{l_{1},a_{1}}E[Y_{2}-Y_{1}|A_{0}=g(a_{0},l_{0}),A_{1}=g(l_{0},l_{1},g(l_{0},a_{0}),a_{1}),L_{0}=l_{0},L_{1}=l_{1}]p(l_{1},a_{1}|A_{0}=g(a_{0},l_{0}),L_{0}=l_{0})da_{1}dl_{1}$
$=E[\mu_{12}^{d}(l_{0},L_{1},g(a_{0},l_{0}),g(l_{0},L_{1},g(l_{0},a_{0}),A_{1})|A_{0}=g(l_{0},a_{0}),L_{0}=l_{0}]$
$\equiv\mu_{02}^{d}(g(l_{0},a_{0}),l_{0})$
where $\mu_{12}^{d}(l_{0},l_{1},a_{0},a_{1})=E[Y_{2}-Y_{1}|L_{0}=l_{0},L_{1}=l_{1},A_{0}=a_{0},A_{1}=a_{1}]$. The derivation is simply repeated applications of iterated expectations, parallel trends, and consistency. It follows that for parametric model $(\gamma_{01}^{g}(l_{0},a_{0}),\gamma_{12}^{g}(h_{1},a_{1}),\gamma_{02}^{g}(l_{0},a_{0}))=(\gamma_{01}^{g}(l_{0},a_{0};\psi^{*}),\gamma_{12}^{g}(h_{1},a_{1};\psi^{*}),\gamma_{02}^{g}(l_{0},a_{0};\psi^{*}))$, parameter $\psi^{*}$ satisfies
$E\begin{bmatrix}q_{01}(L_{0},A_{0})\{Y_{1}-Y_{0}-\gamma_{01}^{g}(L_{0},A_{0};\psi^{*})-\mu^{d}_{01}(L_{0},g(L_{0},A_{0}))\}\\ q_{12}(H_{1},A_{1})\{Y_{2}-Y_{1}-\gamma_{12}^{g}(H_{1},A_{1};\psi^{*})-\mu^{d}_{12}(H_{1},g(H_{1},A_{1}))\}\\ q_{02}(L_{0},A_{0})\{Y_{2}-Y_{1}-\gamma_{12}^{g}(H_{1},A_{1};\psi^{*})-\gamma_{02}^{g}(L_{0},A_{0};\psi^{*})+\gamma_{01}^{g}(L_{0},A_{0};\psi^{*})-\mu_{02}^{d}(L_{0},g(L_{0},A_{0}))\}\end{bmatrix}=0,$ (20)
which again leads to natural estimating equations. In particular, a consistent and asymptotically normal estimator of $\psi^{*}$ can be obtained by solving
$\mathbb{P}_{N}\begin{bmatrix}b_{01}(L_{0},A_{0})\{Y_{1}-Y_{0}-\mu^{d}_{01}(L_{0},A_{0};\beta)\\ b_{12}(\bar{L}_{1},\bar{A}_{1})\{Y_{2}-Y_{1}-\mu^{d}_{12}(\bar{L}_{1},\bar{A}_{1};\beta)\\ b_{02}(L_{0},A_{0})\{\mu^{d}_{12}(L_{0},L_{1},g(L_{0},A_{0}),g(L_{0},L_{1},g(L_{0},A_{0}),A_{1}))-\mu^{d}_{02}(L_{0},A_{0};\beta)\}\\ q_{01}(L_{0},A_{0})\{Y_{1}-Y_{0}-\gamma_{01}^{g}(L_{0},A_{0};\psi)-\mu^{d}_{01}(L_{0},g(L_{0},A_{0}))\}\\ q_{12}(H_{1},A_{1})\{Y_{2}-Y_{1}-\gamma_{12}^{g}(H_{1},A_{1};\psi)-\mu^{d}_{12}(H_{1},g(H_{1},A_{1}))\}\\ q_{02}(L_{0},A_{0})\{Y_{2}-Y_{1}-\gamma_{12}^{g}(H_{1},A_{1};\psi)-\gamma_{02}^{g}(L_{0},A_{0};\psi)+\gamma_{01}^{g}(L_{0},A_{0};\psi)-\mu_{02}^{d}(L_{0},g(L_{0},A_{0}))\}\end{bmatrix}=0$ (21)
where $\mu_{mk}^{d}(\bar{A}_{m},\bar{L}_{m};\beta)$ are models for $\mu_{mk}^{d}(\bar{A}_{m},\bar{L}_{m})$ smooth in an $r$-dimensional parameter $\beta$ (equal to $\beta^{*}$ under the true law) and $q_{mk}(\bar{A}_{m},\bar{L}_{m})$ and $b_{mk}(\bar{A}_{m},\bar{L}_{m})$ are conformable analyst selected index functions. Derivation of Neyman orthogonal estimators is again left for future work.
5 Simulations
5.1 Point exposure
We simulated data according to the data generating process:
$L\sim N(0,1)$
$A|L\sim N(m(L),1)\text{ with }m(L)=\theta_{0}+\theta_{1}L+\theta_{2}L^{2}$
$Y=\xi_{0}+\xi_{1}L+\xi_{2}L^{2}+(\beta_{0}+\beta_{1}L)A+\epsilon;\text{ }\epsilon\sim N(0,1)$
We then sought to estimate the heterogeneous effects of the shift MTP $g(A,L)=A+\delta$. Are interested in the SNMM
$\gamma(A,L)=E[Y|A,L]-E[Y|A+\delta,L]=-\delta(\beta_{0}+\beta_{1})L$
Thus, we specify the correct parametric SNMM
$\gamma(A,L;\psi)=-\delta s(L)^{T}\psi$
with $s(L)=(1,L)$ and $\psi=(\beta_{0},\beta 1)$. We then define
$\tilde{q}(a,l)=q(a-\delta,l)\frac{p(a-\delta|l)}{p(a|l)}$
and choose $q(A,L)=s(L)$. We set true parameter values to be:
$\theta=(0.2,0.8,-0.4);\text{ }\xi=(0.3,0.5,0.2);\text{ }\beta=(0.8,-0.6).$
We then estimated the effect of a shift by $\delta=0.5$ by solving the estimating equations of Theorem 1 with cross-fitting for data sets of sizes 400, 1000, and 3000. We used correctly specified parametric nuisance models. Table 1 shows that the estimator is unbiased and that sandwich confidence intervals had nominal coverage for each component of $\psi$.
Table 1: Monte Carlo performance — point-exposure continuous-shift orthogonal estimator (stable score; $R=200$).
$\psi_{0}$ $\psi_{1}$
n Bias RMSE EmpSD Mean SE Cov95 Bias RMSE EmpSD Mean SE Cov95
$400$ $0.002$ $0.056$ $0.056$ $0.057$ $0.96$ $0.004$ $0.065$ $0.065$ $0.059$ $0.93$
$1000$ $0.001$ $0.037$ $0.037$ $0.035$ $0.93$ $0.002$ $0.037$ $0.037$ $0.035$ $0.93$
$3000$ $0.000$ $0.020$ $0.020$ $0.020$ $0.94$ 0.002-0.002 $0.020$ $0.019$ $0.020$ $0.96$
Notes: Bias = mean of $\hat{\psi}-\psi^{\star}$; EmpSD = empirical SD of $\hat{\psi}$ across replications; Mean SE = average sandwich SE; Cov95 = empirical coverage of 95% Wald CIs from sandwich SEs.
5.2 Time-varying treatment
We consider a two–time-point longitudinal setting with baseline covariate $L_{0}$, treatments $A_{0},A_{1}\in\mathbb{R}$, intermediate covariate $L_{1}$, and outcome $Y\in\mathbb{R}$. The modified treatment policy (MTP) shifts each treatment by a constant
$g_{0}(a_{0})=a_{0}+\delta_{0},\qquad g_{1}(a_{1})=a_{1}+\delta_{1},$
with fixed $\delta_{0},\delta_{1}>0$. The blip functions relative to $g=(g_{0},g_{1})$ are parameterized by low–dimensional bases
$s_{0}(H_{0})=(1,L_{0})^{\top},\qquad s_{1}(H_{1})=(1,L_{1})^{\top},$
and true parameters $\psi_{0}=(\psi_{0,0},\psi_{0,1})^{\top}$, $\psi_{1}=(\psi_{1,0},\psi_{1,1})^{\top}$, so that
$\gamma_{0}^{g}(H_{0},A_{0})\;=\;-\delta_{0}\,s_{0}(H_{0})^{\top}\psi_{0},\qquad\gamma_{1}^{g}(H_{1},A_{1})\;=\;-\delta_{1}\,s_{1}(H_{1})^{\top}\psi_{1}.$
Observed data–generating process (DGP).
We calibrated our generating process such that we could derive the true parameter values for our blip model. Let $n$ i.i.d. observations be generated as follows; all noises are mutually independent and independent of past history.
$L_{0}$ $\sim\mathcal{N}(0,1),$
$A_{0}\mid L_{0}$ $\sim\mathcal{N}\!\big(\,0.4\,L_{0},\;\sigma_{A_{0}}^{2}\,\big),$
$L_{1}\mid(L_{0},A_{0})$ $:=\rho_{0}+\rho_{1}L_{0}+\rho_{2}A_{0}+\nu,\qquad\nu\sim\mathcal{N}(0,\sigma_{L_{1}}^{2}),$
$A_{1}\mid(L_{0},L_{1},A_{0})$ $\sim\mathcal{N}\!\big(\,\kappa_{0}+\kappa_{2}L_{0},\;\sigma_{A_{1}}^{2}\,\big),$
$\mu_{1}(H_{1},a_{1})$ $:=b_{1}(L_{0},L_{1})+\{\psi_{1,0}+\psi_{1,1}L_{1}\}\,a_{1},$
$Y\mid(H_{1},A_{1})$ $:=\mu_{1}(H_{1},A_{1})+\varepsilon,\qquad\varepsilon\sim\mathcal{N}(0,\sigma_{Y}^{2}),$
where $b_{1}(L_{0},L_{1})=\beta_{10}+\beta_{1L_{1}}L_{1}+\beta_{1L_{0}}L_{0}$ is a baseline outcome component. With this construction,
$\mu_{1}(H_{1},a_{1})-\mu_{1}\!\big(H_{1},a_{1}+\delta_{1}\big)\;=\;-\delta_{1}\{\psi_{1,0}+\psi_{1,1}L_{1}\}\;=\;\gamma_{1}^{g}(H_{1},A_{1})$
holds by design for any $H_{1}$.
To ensure the time-0 blip identity also holds,
$\mu_{0}(H_{0},a_{0})-\mu_{0}\!\big(H_{0},a_{0}+\delta_{0}\big)\;=\;-\delta_{0}\,s_{0}(H_{0})^{\top}\psi_{0}\;=\;\gamma_{0}^{g}(H_{0},A_{0}),$
we calibrate $\psi_{0}$ so that the slope in $a_{0}$ of
$\mathbb{E}\!\left[\,\mu_{1}\!\big(H_{1},A_{1}+\delta_{1}\big)\,\middle|\,H_{0},\ A_{0}=a_{0}\right]$
equals $s_{0}(H_{0})^{\top}\psi_{0}$ for all $L_{0}$. Under the simple $A_{1}$ law above (mean depending on $L_{0}$ but not on $A_{0}$ or $L_{1}$), a short calculation yields
$\psi_{0,0}\;=\;\rho_{2}\,\beta_{1L_{1}}\;+\;\rho_{2}\,\psi_{1,1}\,(\kappa_{0}+\delta_{1}),\qquad\psi_{0,1}\;=\;\rho_{2}\,\psi_{1,1}\,\kappa_{2}$
so that $\mu_{0}(H_{0},a_{0}):=\mathbb{E}\!\left[\,\mu_{1}\!\big(H_{1},A_{1}+\delta_{1}\big)\,\middle|\,H_{0},\ A_{0}=a_{0}\right]$ is linear in $a_{0}$ with slope $s_{0}(H_{0})^{\top}\psi_{0}$.
Default parameter values.
In all experiments we use
$\delta_{0}=0.4,\quad\delta_{1}=0.5,\quad\psi_{1}=(0.5,\,0.3),\quad(\rho_{0},\rho_{1},\rho_{2})=(0.1,\,0.6,\,0.8),$
$(\kappa_{0},\kappa_{2})=(0.2,\,0.35),\quad(\beta_{10},\beta_{1L_{1}},\beta_{1L_{0}})=(0.25,\,0.5,\,0.2),$
and standard deviations $(\sigma_{A_{0}},\sigma_{L_{1}},\sigma_{A_{1}},\sigma_{Y})=(1.0,\,0.5,\,1.0,\,1.0)$. The calibration then implies the true time-0 blip coefficients
$\psi_{0,0}=0.8\times 0.5+0.8\times 0.3\times(0.2+0.5)=0.568,\qquad\psi_{0,1}=0.8\times 0.3\times 0.35=0.084.$
Thus the overall target parameter vector is
$\psi^{\star}=\big(\psi_{0,0},\psi_{0,1},\psi_{1,0},\psi_{1,1}\big)^{\top}=\big(0.568,\ 0.084,\ 0.5,\ 0.3\big)^{\top}.$
Results
Table 2 displays the bias, bootstrap coverage, and empirical standard errors over 500 data sets of size n=1,000. We observe very low bias and near nominal coverage.
Table 2: Monte Carlo (R=500) for longitudinal MTP estimator with continuous shifts; bootstrap 95% coverage reported.
$\psi_{0,0}$ $\psi_{0,1}$ $\psi_{1,0}$ $\psi_{1,1}$
n Bias RMSE EmpSD Bias RMSE EmpSD Bias RMSE EmpSD Bias RMSE EmpSD
$1000$ 0.000-0.000 $0.049$ $0.049$ 0.002-0.002 $0.039$ $0.039$ $0.000$ $0.033$ $0.033$ $0.002$ $0.026$ $0.026$
Boot 95% cov. 0.93 0.94 0.96 0.96
6 Real Data Application: Effect of Shifting Workplace Mobility on Covid Incidence
Data.
We linked the Google Community Mobility Reports [Google LLC, 2022] (workplaces index; daily % point deviation from the Jan 3–Feb 6, 2020 baseline) with New York Times county-level COVID-19 case counts [The New York Times, 2025]. Units were counties. The exposure $A$ is the 7-day average of the workplace mobility index ending at $t_{0}$ (units: percentage points change from baseline mobility in the January 3- February 6 period). We defined the calendar date $t_{0}$ to be the date with a mobility measurement closest to June 1, 2020, breaking ties arbitrarily. We excluded 74 counties without an eligible $t_{0}$. Outcomes are future incident cases per 100,000 over the 7 days beginning 14 days after $t_{0}$.
Intervention (MTP).
We consider a constant shift policy $g(a)=a+\delta$ with $\delta=-5$ percentage points. This means that under this intervention each county’s percentage point change in mobility score from Jan3-Feb6 would be shifted down 5 percentage points from its observed value.
Adjustment covariates.
We adjusted for: population, Rural–Urban Continuum (RUC) score (1=most urban, 9=most rural), percent Black, percent Hispanic, Republican vote share in 2016, poverty rate, unemployment rate, uninsurance rate, land area, population density, and per-capita income.
Estimand and blip specification.
We model the blip as linear in the RUC score:
$\Delta(H)\equiv\mathbb{E}\{Y(A+\delta)-Y(A)\mid H\}\approx\delta\,s(H)^{\top}\psi,\qquad s(H)=(1,\ \mathrm{RUC})^{\top}.$
Here, $\psi_{1}$ is the average effect per 1 p.p. shift at RUC=0=0 (intercept on the RUC scale) and $\psi_{2}$ captures linear effect modification by RUC. With $\delta=-5$, a positive $\psi_{2}$ means that a decreasing mobility is more protective in more rural areas. Treating RUC as a nominal variable when it is really ordinal is a slight abuse. It surely pales in comparison to the omission of other covariates from the blip function. Keep in mind that this is an illustrative analysis.
Estimator.
[…]
Results.
Estimated blip coefficients (per 1 p.p. shift):
$\hat{\psi}_{1}=14.3\quad(\mathrm{SE}=14.0;\ 95\%\ \mathrm{CI}:\ -13.2,\ 41.8),\qquad\hat{\psi}_{2}=-3.1\quad(\mathrm{SE}=5.0;\ 95\%\ \mathrm{CI}:\ -13.0,\ 6.8).$
With the policy shift $\delta=-5$ p.p., the implied change in cases per 100,000 at a given RUC is $\widehat{\Delta}(\mathrm{RUC})=\delta\,(\hat{\psi}_{1}+\hat{\psi}_{2}\,\mathrm{RUC})$. Delta-method estimates (and 95% CIs) at representative RUC values are:
$\begin{array}[]{lccc}\text{RUC}=1:&\widehat{\Delta}=-7.24,\ \mathrm{SE}=38.08,\ \mathrm{CI}\ [-81.88,\ 67.40];\\ \text{RUC}=5:&\widehat{\Delta}=-20.79,\ \mathrm{SE}=22.46,\ \mathrm{CI}\ [-64.82,\ 23.24];\\ \text{RUC}=9:&\widehat{\Delta}=-34.34,\ \mathrm{SE}=41.65,\ \mathrm{CI}\ [-115.97,\ 47.29].\end{array}$
Interpretation.
Point estimates suggest that a 5 p.p. reduction in workplace mobility index relative to what was observed would have lowered future COVID-19 cases per 100,000, with larger reductions in more rural counties (more negative at higher RUC). However, all 95% CIs include zero, and effects are imprecisely estimated in this purely illustrative analysis.
7 Discussion and Future Work
Estimating heterogeneous effects of MTPs with SNMMs can be both of scientific interest and important for planning realistic interventions. Suppose policy makers want to know the effect of a campaign to decrease opioid prescription dosing on some continuous quality of life utility measure. They believe the impact of the program might be approximated by the effect of an MTP: ‘prescribe a 20% lower dose of opioids than you normally would’. One might be interested in how the effect of this intervention varies with the natural dose. Perhaps it turns out that reductions are more impactful when the natural dose is moderate, as opposed to low or high. Learning this might lead to hypotheses about the mechanism of opioid addiction that could be tested in further studies. The knowledge might also help to design targeted opioid reduction interventions.
Furthermore, economists who commonly apply DiD can now study MTPs (albeit only using parametric nuisance models for the time being) under similar parallel trends assumptions. For example, they can estimate the effect of increasing minimum wage by 1 dollar more than it actually increased, and how this effect varies with the amount it actually increased. This is not an estimand that can be targeted by standard DiD.
There are many directions for future work. First, it remains to derive Neyman orthogonal estimators under parallel trends assumptions. Given past modifications of SNMM results to accommodate parallel trends [Shahn et al., 2022], we believe this should be achievable. Second, SNMMs can also be estimated under instrumental variable assumptions [Robins, 1994b], and instrumental variable estimation of MTPs would also be of interest. Third, SNMM variants have been developed to estimate effects on survival [Picciotto et al., 2012] and binary [Wang et al., 2023] outcomes. Effects of MTPs on survival and binary outcomes are of course of interest as well. Fourth, sensitivity analysis for SNMMs [Robins, 2004, Robins et al., 2000] is well developed and would hopefully easily port over to the MTP setting, but that should be confirmed. Finally, marginal SNMMs that model heterogeneity as a function of a subset of covariates required for adjustment have been developed elsewhere, and it might be useful to transport them to the MTP setting, too.
8 Acknowledgments
I am extremely indebted to Jamie Robins for lots of things, but most relevant for this Acknowledgments section is that he suggested the general idea for this paper (‘SNMMs for MTPs’). However, all errors are my own.
Appendix
8.1 Proof of Theorem 1
Proof.
Step 1: Identification. By the SNMM restriction,
$E\!\left[q\{\mu(A,L)-\mu(g(A,L),L)-\gamma_{g}(L,A;\psi^{\ast})\}\right]=0.$
Moreover $E[Y-\mu(A,L)\mid A,L]=0$, hence the augmentation term in (11) has mean zero. Summing yields $E[\phi(O;\psi^{\ast},\eta^{\ast})]=0$.
Step 2: Orthogonality w.r.t. $\mu$. Consider $\mu_{t}=\mu^{*}+t\,\delta\mu+o(t)$ with $\pi$ fixed. Only $Y-\mu_{t}$ and the identifying bracket depend on $t$. Differentiate the expectation of (11) at $t=0$:
$\left.\frac{d}{dt}\right|_{0}E\!\left[(q-\tilde{q})(Y-\mu_{t})\right]=-\,E\!\left[(q-\tilde{q})\delta\mu\right]=-\,\langle q-\tilde{q},\delta\mu\rangle_{w}.$
For the identifying term,
$\left.\frac{d}{dt}\right|_{0}E\!\left[q\{\mu_{t}-T_{g}\mu_{t}-\gamma_{g}(\psi^{*})\}\right]=E\!\left[q\,\delta\mu\right]-E\!\left[q\,T_{g}\delta\mu\right].$
Rewrite under $\langle\cdot,\cdot\rangle_{w}$ and apply the definition of $\tilde{q}$:
$E[q\,\delta\mu]-E[q\,T_{g}\delta\mu]=\langle q,\delta\mu\rangle_{w}-\langle q,T_{g}\delta\mu\rangle_{w}=\langle q,\delta\mu\rangle_{w}-\langle\tilde{q},\delta\mu\rangle_{w}=\langle q-\tilde{q},\delta\mu\rangle_{w}.$
Summing the two derivatives: $\langle q-\tilde{q},\delta\mu\rangle_{w}+\langle q-\tilde{q},\delta\mu\rangle_{w}=0$.
Step 3: Orthogonality w.r.t. $\pi$. Let $\pi_{t}=\pi^{*}+t\,\delta\pi+o(t)$ and $w_{t}=\pi_{t}$. Then both $w_{t}$ and $\tilde{q}_{t}:=T_{g}^{\dagger}q$ vary with $t$. Write
$F(t):=E\!\left[\phi(O;\psi^{*},\mu^{*},\pi_{t})\right]=\underbrace{E\!\left[(q-\tilde{q}_{t})(Y-\mu^{*})\,\right]}_{=:A(t)}+\underbrace{E\!\left[q\{\mu^{*}-T_{g}\mu^{*}-\gamma_{g}(\psi^{*})\}\right]}_{=:B\ \text{(const.\ in $t$)}}.$
Hence $F^{\prime}(0)=A^{\prime}(0)$. Differentiate $A(t)$ at 0:
$A^{\prime}(0)=-\,E\!\big[\dot{\tilde{q}}\,(Y-\mu^{*})\,w^{*}\big]\;+\;E\!\big[(q-\tilde{q}^{*})(Y-\mu^{*})\,\dot{w}\big],$
where dots denote $t$-derivatives at 0, and $w^{*}=\pi^{*}$. Now condition on $(A,L)$. Since $E[Y-\mu^{*}(A,L)\mid A,L]=0$, both terms vanish:
$E\!\big[\dot{\tilde{q}}\,(Y-\mu^{*})\,w^{*}\big]=E\!\big[\,E[Y-\mu^{*}\mid A,L]\;\dot{\tilde{q}}\,w^{*}\,\big]=0,$
and similarly for the second term. Therefore $A^{\prime}(0)=0$, i.e. $F^{\prime}(0)=0$.
An equivalent route is to differentiate the identity $\langle q,T_{g}h\rangle_{w_{t}}=\langle\tilde{q}_{t},h\rangle_{w_{t}}$ at $t=0$ with $h(a,l)=Y-\mu^{*}(a,l)$. This yields
$E\!\left[(q-\tilde{q}^{*})\,T_{g}h\,\dot{w}\right]=E\!\left[\dot{\tilde{q}}\,h\,w^{*}\right],$
which is exactly the cancellation needed in $A^{\prime}(0)$ when $h$ is replaced by $Y-\mu^{*}$. ∎
8.2 Proof of Theorem 2
Proof.
Step 1: Identification. By construction, for each $t$, $E\!\big[V_{t+1}-\mu_{t}(H_{t},A_{t})\mid H_{t},A_{t}\big]=0,$ hence the augmentation term of $\phi_{t}$ has mean zero. The longitudinal SNMM restriction gives $E\!\big[q_{t}\{\mu_{t}(H_{t},A_{t})-\mu_{t}(H_{t},g_{t}(A_{t},H_{t}))-\gamma_{t}(H_{t},A_{t};\psi^{*})\}\big]=0.$ Summing over $t$ yields $E[\Phi(O;\psi^{*},\eta^{*})]=0$.
Step 2: Orthogonality w.r.t. $\mu_{t}$. Fix $t$ and perturb $\mu_{t,\varepsilon}=\mu_{t}^{*}+\varepsilon\delta\mu_{t}+o(\varepsilon)$, holding all $\{\mu_{s},\pi_{s}\}_{s\neq t}$ and $\pi_{t}$ fixed. Note that $V_{t+1}$ is treated as fixed when differentiating $\phi_{t}$ in its own nuisances. Differentiate the expectation of $\phi_{t}$ at $\varepsilon=0$:
$\left.\frac{d}{d\varepsilon}\right|_{0}E\!\left[(q_{t}-\tilde{q}_{t})(V_{t+1}-\mu_{t,\varepsilon}(H_{t},A_{t}))\right]=-\,\langle q_{t}-\tilde{q}_{t},\delta\mu_{t}\rangle_{w_{t}},$
and
$\left.\frac{d}{d\varepsilon}\right|_{0}E\!\left[q_{t}\{\mu_{t,\varepsilon}-T_{g,t}\mu_{t,\varepsilon}-\gamma_{t}(\psi^{*})\}\right]=\langle q_{t},\delta\mu_{t}\rangle_{w_{t}}-\langle q_{t},T_{g,t}\delta\mu_{t}\rangle_{w_{t}}=\langle q_{t}-\tilde{q}_{t},\delta\mu_{t}\rangle_{w_{t}},$
by the adjoint identity for time $t$. The two derivatives cancel exactly. Because $\phi_{t}$ does not depend on $\mu_{s}$ for $s\neq t$ (given $V_{t+1}$ fixed), we obtain $D_{\mu_{t}}E[\phi_{t}]=0$.
Step 3: Orthogonality w.r.t. $\pi_{t}$. Perturb $\pi_{t,\varepsilon}$; then $w_{t,\varepsilon}=\pi_{t,\varepsilon}$ and $\tilde{q}_{t,\varepsilon}=T_{g,t}^{\dagger,w_{t,\varepsilon}}q_{t}$ vary with $\varepsilon$. Write
$F_{t}(\varepsilon):=E[\phi_{t}(O;\psi^{*},\mu_{t}^{*},\pi_{t,\varepsilon})]=\underbrace{E\!\left[(q_{t}-\tilde{q}_{t,\varepsilon})\,(V_{t+1}-\mu_{t}^{*})\,\right]}_{=:A_{t}(\varepsilon)}+\underbrace{E\!\left[q_{t}\{\mu_{t}^{*}-T_{g,t}\mu_{t}^{*}-\gamma_{t}(\psi^{*})\}\right]}_{=:B_{t}\ \text{const.}}.$
Hence $F_{t}^{\prime}(0)=A_{t}^{\prime}(0)$. Differentiate at 0:
$A_{t}^{\prime}(0)=-\,E\!\big[\dot{\tilde{q}}_{t}\,(V_{t+1}-\mu_{t}^{*})\,w_{t}^{*}\big]\;+\;E\!\big[(q_{t}-\tilde{q}_{t}^{*})\,(V_{t+1}-\mu_{t}^{*})\,\dot{w}_{t}\big].$
Conditioning on $(H_{t},A_{t})$ and using $E[V_{t+1}-\mu_{t}^{*}(H_{t},A_{t})\mid H_{t},A_{t}]=0$, both terms are zero, hence $A_{t}^{\prime}(0)=0$. Equivalently, differentiate the time-$t$ adjoint identity $\langle q_{t},T_{g,t}h\rangle_{w_{t,\varepsilon}}=\langle\tilde{q}_{t,\varepsilon},h\rangle_{w_{t,\varepsilon}}$ at $\varepsilon=0$ with $h(H_{t},A_{t})=V_{t+1}-\mu_{t}^{*}(H_{t},A_{t})$ to see the same cancellation.
Summing the time-local derivatives over $t$ gives orthogonality for $\Phi$. ∎